[figure: 9.7 cm]

MAGNETOELASTIC AND OPTICAL STUDIES OF LOW TEMPERATURE MAGNETIC RELAXATION IN RARE-EARTH

ORTHOALUMINATES WITH ISING IONS

I. Krynetskii^{a}, A. Popkov^{b}, M. Borowiec^{c} and H. Szymczak^{c} aMoscow State University, 119899, Moscow, Russia

bState Research Institute for Physical Problems, 103460, Zelenograd, Moscow, Russia

cInstitute of Physics, Polish Academy of Science, Al. Lotników, 02-668, Warsaw, Poland

The new class of magnetic materials with low-temperature magnetic relaxation was studied. The volume crystal DyAlO_{3} with antiferromagnetic ordered high anisotropy rare-earth ions had the exponential magnetic relaxation was found. For the first time the low temperature relaxation was investigated by means of magnetostriction and rareearth absorption spectra instead of a routine magnetization measurements. The rareearth subsystem metastable state was created as a result of a thermal shock at the sample temperature passing through T =2.18 K. The shift of DyAlO_{3} absorption spectra for metastable and equilibrium states was evaluated as 10^{−1}. The possible mesoscopic nature of tunneling process is under consideration. This work was partly supported by RFBR (grant 98-02-16469) and Russian Education Ministry (grant 97-0-7.3-157).

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : I.Krynetskii

Address for correspondence : Department of Physics Moscow State University 119899, Moscow, Russia

Email address : krynets@plms.phys.msu.su